1.) (r^-4) ( -6r^2t^2)
2.) w^3 / -6w^-6
3.) (k^6) (k^3)^5
4.) t^-9 / g^-9

Answer by RAY100(1637)

Remember negative exponent is reciprocal,,,,a^-2 = 1/a^2
.
AND , to divide common bases,,,substract exponents,,,,a^4/a^2=a^(4-2) = a^2
.
(1) (r^-4)(-6r^2t^2) = -6 r^-2t^2= -6t^2 / r^2
.
(2) w^3 / (-6w^-6) = w^9/-6 = (-1/6) w^9
.
(3) k^6 / (k^3)^5 = k^6/k^15 =k^-9 = 1/k^9
.
(4) t^-9 / q^-9 = q^9/t^9 = (q/t)^9
.
check by assigning numbers to variables and checking original and answer,,,all ok